OpenStudy (anonymous):

Help! The slope is -3, and the line passes through the point (4,5). The equation is: a. 3x-y+1=0 b. 3x+y-17=0 c. 3x+y-19=0 d. -3x+y-7=0

OpenStudy (anonymous):

slope = -3 (y-5)=3(x-4) (y-5)-3(x-4)=0 You will then get the ans by simplification

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you are welcome